OpenStudy (anonymous):

During his hockey career, Gordie Howe played 1,767 NHL regular season games, scoring 801 goals and making 1,049 assists. If goals and assists each count as one point, what was his scoring rate in points per game? Round to the nearest hundredth. *

OpenStudy (anonymous):

HOCKEY OH YA

OpenStudy (anonymous):

1049+801= X 1767/X= Do you know how to do that? Do you know why I am doing that? @richytyb

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

We need to figure out the point number. That's why we do this. 1049+801= X Solve for X!